# Leveraging Moving Sound Source Trajectories for Universal Sound Separation

### v1

@@ IV-B2 Ablation study
Additionally, we plot the angular error between the estimated trajectories and the target trajectories under different amplitudes of the energy envelope. Fig 7 shows the results. It can be observed that when the sound source energy is low, the angular error of ENtrack and Mixtrack is nearly the same. This indicates that locating the sound source becomes challenging when its energy is low. Regardless of the weights assigned to moments of low energy, the network can hardly learn from these moments. Due to ENtrack assigning smaller weights to the trajectory loss for low-energy moments, the differential loss carries a larger proportion in the overall loss. As a result, the network can learn more from the differential loss, which contains the motion information of the sound source. This allows the network to utilize contextual information to determine the position of the sound source at that moment. Thus the angular error of ENtrack is slightly lower than Mixtrack. As the amplitude of the energy envelope increases, the advantage of ENtrack over Mixtrack becomes increasingly evident, which is because that ENtrack applies more weights to these high-energy moments than low-energy moments during one signal period thus the network can learn more from high-energy moments. Tracking the sound source based on sound source energy offers the greatest potential to improve the performance during periods of high sound source activity.
@@
TABLE II: The EWRMSAE of trajectories from different rounds of the mutual facilitation.
@@
init-track 18.97
MFTS-track-iter1 14.89
MFTS-track-iter2 14.62
MFTS-track-iter3 14.62
To show the efficacy of mutual facilitation mechanism, we perform the mutual facilitation three times and compare the results. Firstly, we compare the EWRMSAE values for initial tracking, the first time precise tracking, the second time precise tracking, and the third time precise tracking, represented as init-track, MFTS-track-iter1, MFTS-track-iter2, and MFTS-track-iter3, respectively. The result is shown in Table II. It can be observed that performing precise tracking on the results of multichannel separation significantly improves sound source tracking performance. By extracting multi-channel signals based on this precise tracking and then do a further precise tracking, the tracking performance is further enhanced. Additionally, as shown in Table II, after two rounds of mutual facilitation, the tracking performance reaches its optimal value, eliminating the need for further mutual facilitation.
@@
TABLE III: The separation performance from different rounds of the mutual facilitation and from the baseline.
Method SI-SNR SNR SDR
SpatialNet-mc 12.26 12.98 13.17
MFTS-iter1 13.12 15.37 14.90
MFTS-iter2 13.59 15.74 15.27
MFTS-iter3 13.62 15.76 15.30
Additionally, we compare the performance of target sound extraction based on initial tracking, the first time precise tracking after one round of the mutual facilitation and the second time precise tracking after two rounds of the mutual facilitation. These three methods are represented as MFTS-iter1, MFTS-iter2, MFTS-iter3. The SpatialNet directly separates the multi-channel signals without trajectory inputs is also compared, which is denoted as SpatialNet-mc. The result is shown in Table III. It can be observed that, compared to target sound extraction based on initial tracking, the separation performance based on the precise tracking is significantly enhanced and markedly superior to the baseline method. Moreover, it can be seen that after one round of mutual facilitation, the separation results obtained in the second round of mutual facilitation reach near the optimal value. This is because the precise tracking performance approaches its optimal value, and the improvement is not significant after the second round of mutual facilitation. By analyzing the results from Table II and Table III, we choose to perform two rounds of mutual facilitation and select the separation and tracking results from the second round for the sake of effectiveness and computational simplicity.
@@
TABLE IV: The separation performance of channel 00 before and after neural beamforming.
Method SI-SNR SNR SDR
MFTS-WOBF 14.33 15.27 15.74
MFTS 15.21 16.07 16.55
@@
Table IV shows the separation performance on channel 0 for the multichannel separation results obtained in the second round of mutual facilitation before neural beamforming (MFTS-WOBF), as well as the separation performance after neural beamforming (MFTS). It can be observed that neural beamforming utilizes the separated multichannel information to further enhance the separation results on the reference channel.
@@ V Conclusion
In this paper, we propose a method that leverages the mutual facilitation mechanism between sound source separation and tracking to generate precise trajectories of moving sound sources for universal sound separation. Initially, we estimate the sound energy envelope of each source, which serves as the basis for tracking and deriving their trajectories. Following this, target sound extraction is performed using the obtained trajectories, allowing for more precise sound source tracking based on the separated signals. The improved tracking accuracy further enhances sound separation, creating a loop that can be repeated for increasingly accurate separation and trajectory results. Experimental results demonstrate that our proposed method outperforms baseline approaches, achieving superior performance based on the precise trajectories enabled by this mutual facilitation mechanism. This approach underscores the effectiveness of integrating sound source localization and separation to mutually enhance their respective performances.

### v2

@@ IV-C1 Initial tracking
Additionally, we statistically analyze the envelope signal value and the angular error at every time sample, plotting the mean angular errors across different envelope signal magnitude ranges. Fig. 11 shows the results. It can be observed that when the sound source energy is low, init-track and mix-track demonstrate closely comparable error levels. This indicates that locating the sound source becomes challenging when its energy is low. Regardless of the weights assigned to moments of low energy, the network can hardly learn from these moments. Due to init-track assigning smaller weights to the trajectory loss for low-energy moments, the differential loss carries a larger proportion in the overall loss. As a result, the network can learn more from the differential loss, which contains the motion information of the sound source. This allows the network to utilize contextual information to determine the position of the sound source at that moment. Thus the angular error of init-track is lower than mix-track. As the amplitude of the envelope signal increases, the advantage of init-track over mix-track becomes increasingly evident, which is because that init-track applies more weights to these high-energy moments than low-energy moments during one signal period thus the network can learn more from high-energy moments. Tracking the sound source based on sound source energy offers the greatest potential to improve the performance during periods of high sound source activity.
@@
TABLE III: The EWRMSAE () of trajectories from different rounds of the mutual facilitation.
@@
init-track 19.10
MFTS-track-iter1 14.66
MFTS-track-iter2 14.30
MFTS-track-iter3 14.30
@@
TABLE IV: The separation performance from different rounds of the mutual facilitation and from the baselines.
Method SNR SI-SNR SDR
MC-SpatialNet 12.04 11.15 12.60
Env-only 13.41 12.62 13.94
MFTS-iter1 13.86 12.72 14.21
MFTS-iter2 14.12 13.09 14.51
MFTS-iter3 14.14 13.12 14.53
@@
TABLE V: The separation performance of channel 00 before and after neural beamforming.
Method SNR SI-SNR SDR
MFTS-WOBF 14.87 13.92 15.37
MFTS 15.35 14.48 15.83
IV-C2 Mutual facilitation
To show the efficacy of mutual facilitation mechanism, we perform the mutual facilitation three times and compare the results. Firstly, we compare the EWRMSAE values for initial tracking, the first time precise tracking, the second time precise tracking, and the third time precise tracking, represented as init-track, MFTS-track-iter1, MFTS-track-iter2, and MFTS-track-iter3, respectively. The result is shown in Table III. It can be observed that performing precise tracking on the results of multi-channel separation significantly improves sound source tracking performance. By extracting multi-channel signals based on this precise tracking and then do a further precise tracking, the tracking performance is further enhanced. Additionally, as shown in Table III, after two rounds of mutual facilitation, the tracking performance reaches its optimal value, eliminating the need for further mutual facilitation.
Additionally, we compare the performance of target sound extraction based on the results of initial tracking, the first time precise tracking after one round of the mutual facilitation and the second time precise tracking after two rounds of the mutual facilitation. These three methods are represented as MFTS-iter1, MFTS-iter2, MFTS-iter3. The SpatialNet directly separates the multi-channel signals without trajectory inputs (MC-SpatialNet) is also compared. We also compare the method using only the magnitude of the intensity trajectory (Env-only), which is indeed the envelope signal estimation, against the approach utilizing complete intensity trajectories (both direction and magnitude), to validate the effectiveness of utilizing spatial information for the proposed method. The result is shown in Table IV. It can be observed that, compared to target sound extraction based on initial tracking, the separation performance based on the precise tracking is significantly enhanced and markedly superior to the baseline method. Moreover, it can be seen that after one round of mutual facilitation, the separation results obtained in the second round of mutual facilitation reach near the optimal value. This is because the precise tracking performance approaches its optimal value, and the improvement is not significant after the second round of mutual facilitation. The Env-only method demonstrates inferior performance compared to the full trajectory method, further validating the effectiveness of spatial information utilization for separation in our framework. By analyzing the results from Table III and Table IV, we choose to perform two rounds of mutual facilitation and select the separation and tracking results from the second round for the sake of effectiveness and computational simplicity.
TABLE VI: The NMSE (dB) of envelope estimation and accuracy (%) of source count estimation under different number of sources.
Metrics 2 sources 3 sources
NMSE -17.85 -13.11
accuracy 92.22 84.06
TABLE VII: The EWRMSAE () under different number of sources.
Method 2 sources 3 sources
init-track 19.50 27.34
MFTS-track-iter1 15.09 21.86
MFTS-track-iter2 14.58 21.14
TABLE VIII: The separation performance under different number of sources.
Method 2 sources 3 sources
SNR SI-SNR SDR SNR SI-SNR SDR
MFTS-iter1 12.24 11.20 12.80 7.56 6.03 8.00
MFTS-iter2 12.66 11.82 13.30 8.22 6.70 8.48
MFTS-WOBF 13.35 12.57 14.13 8.64 7.25 9.08
MFTS 13.49 13.00 14.35 8.93 7.83 9.30
IV-C3 Neural beamforming
@@
Table V shows the separation performance on the single-channel zero-th order signal, which is equivalent to the signal recorded by an omnidirectional microphone placed at the center of a spherical microphone array, for the multi-channel separated FOA signals obtained in the second round of mutual facilitation before neural beamforming (MFTS-WOBF, we select the zero-th order to calculate metrics), as well as the separation performance after neural beamforming (MFTS). It can be observed that neural beamforming utilizes the separated multi-channel information to further enhance the separation results on the reference channel.
IV-D Extension to an unknown number of sources
As evident from the preceding descriptions, our method requires prior knowledge of the number of sound sources. This necessity arises from the envelope estimation network’s requirement. The envelope estimation network needs to predefine the number of sound sources to determine its number of output channels. The task of the envelope estimation network is essentially a coarse source separation - not separating source signals themselves, but rather separating the envelopes of source signals. Numerous existing methods have achieved source separation with an unknown number of sources, such as encoder-decoder attractor (EDA) [51],[52], recursion-based methods [53], and approaches constrained on the maximum number of sources [54]. Inspired by [54], We extend the envelope estimation network to handle an unknown number of sources under the premise of knowing the maximum number of sources $C_{max}$. Both the network outputs and training labels have dimensions of $C_{max}\times 1\times T$. For cases where the actual source count $C<C_{max}$, the training labels consist of $C$ ground-truth source envelopes concatenated with $C_{max}-C$ all-zero tensor of dimension $1\times T$. The network training loss for the $c$-th channel is defined as:
$\begin{split}\mathcal{L}_{env,c}&=\begin{cases}\text{NMSELoss}(\mathbf{\hat{E}}^{est}_{c},\mathbf{\hat{E}}^{trg}_{c})&{c<C}\\ \mathcal{L}^{inactive}(\mathbf{\hat{E}}^{est}_{c},\mathbf{\hat{E}}^{trg}_{c})&{C\leq c\leq C_{max}}\end{cases},\\ \end{split}$ (10)
where $\mathcal{L}^{inactive}$ is defined as the loss function when $\mathbf{\hat{E}}^{trg}_{c}$ is an all-zero tensor:
$\begin{split}\mathcal{L}^{inactive}(\mathbf{\hat{E}}^{est}_{c},\mathbf{\hat{E}}^{trg}_{c})=&10\log_{10}(\sum_{t=0}^{T}(\hat{E}^{est}_{c}(t)-\hat{E}^{trg}_{c}(t))^{2}\\ &+\tau\sum_{t=0}^{T}(\hat{E}^{mix}(t))^{2}),\end{split}$ (11)
where $\hat{E}^{trg}_{c}(t)\equiv 0$, $\mathbf{\hat{E}}^{mix}$ denotes the groundtruth envelope of the zero-th order of the audio mixture and $\tau$ equals to $0.01$, which is the same as [54]. The uPIT is also utilized during training. During testing, if the maximum value of an output channel in the envelope estimation network falls below $0.25$, we consider the corresponding channel to contain no active sound source.
To validate the effectiveness of our method under an unknown number of sources, we conducted experiments under an unknown number of sources. The experimental setup mirrored the two-source scenario, except that the number of sources is randomly set to $2$ or $3$ and remained unknown to our method. Table VI presents the NMSE values from the envelope estimation network and the accuracy of source count estimation. The results demonstrate high estimation accuracy for both 2-source and 3-source scenarios. Table VII shows the source tracking performance across different number of sources and mutual facilitation rounds. Table VIII compares the multi-channel separation performance across various mutual facilitation rounds, as well as the single-channel separation results on the zero-th order signal before and after neural beamforming. While a performance degradation is observed under an unknown number of sources, our method still maintains sufficient effectiveness.
@@ V Conclusion
In this paper, we propose a method that leverages the mutual facilitation mechanism between sound source separation and tracking to generate precise trajectories of moving sound sources for universal sound separation. Initially, we estimate the signal envelope of each source, which serves as the basis for tracking, and derive the trajectories of sources. Following this, target sound extraction is performed using the obtained trajectories, allowing for refined sound source tracking based on the separated signals. The improved tracking accuracy further enhances sound separation, creating a loop that can be repeated for increasingly accurate separation and trajectory results. Experimental results demonstrate that our proposed method outperforms baseline approaches, achieving superior performance based on the precise trajectories enabled by this mutual facilitation mechanism. This approach underscores the effectiveness of integrating sound source localization and separation to mutually enhance their respective performances.